# MoveScanner: Analysis of Security Risks of Move Smart Contracts

### v2

@@ 1. Introduction
Blockchain technology has evolved continuously since the birth of Bitcoin, from the introduction of smart contracts by Ethereum, and the emergence of third-generation blockchains such as Solana, Aptos, and Sui in recent years. It has gradually transitioned from a single-purpose value-storage system to a programmable distributed application platform. Smart contracts, as a core component of blockchain, grant the system programmability but also introduce new security risks(luu2016making, ). The Move language, proposed by the Meta team, was designed with security as its core objective. Its innovative resource model treats digital assets as first-class citizens, achieving non-replicability and non-discardability through a linear type system and explicit move operations, thereby preventing resource leaks and double spending at the language level. Combined with a strict type system, formal verification(gong2025information, ), and fine-grained permission controls, Move offers superior security compared to traditional contract languages, such as Ethereum. However, in actual development, Move smart contracts may still contain vulnerabilities due to programming errors, complex business logic, or cross-module interactions(tsankov2018securify, ).
@@
According to Certik’s 2023 Blockchain Security Report, losses caused by smart contract vulnerabilities exceeded \$2.83 billion, with re-entrancy attacks, permission vulnerabilities, and logical errors as the primary attack types. The Ethereum ecosystem has developed a relatively mature security analysis toolchain, including Oyente, Securify, ZEUS, Slither, and Mythril, which utilize symbolic execution, abstract interpretation, or static analysis methods to this study lay the foundation for smart contract security research(li2025scalm, ). In contrast, security tools for the Move ecosystem remain underdeveloped. Move Prover is primarily focused on the formal verification of specific properties, while MoveCheck and VerMove can detect some errors but struggles with complex cross-module vulnerabilities. Manual audit services provided by security companies also lack automation and systematic approaches(tolmach2021survey, ). This highlights the lack of comprehensive, automated security analysis tools for smart contracts on the Move platform. To address this gap, this paper proposes and implements a static analysis tool, MoveScanner. The tool constructs a control flow graph and data flow analysis framework, combined with cross-module call tracing technology, enabling the detection of five critical vulnerability categories, including resource leaks, permission defects, and arithmetic overflows(liu2025sok, ). MoveScanner adopts a modular architecture, supporting bytecode-level analysis and multi-chain adaptation. They introduced resource trace tracking algorithms and capability matrix analysis methods to reduce the number of false positives. Experimental results show that MoveScanner achieves a detection accuracy of 88.2% in benchmark tests, outperforming existing tools in terms of performance(kalra2018zeus, ).
@@ 2.1. Move Language
Move is a smart contract language specifically designed for the secure management and transfer of digital assets. It was proposed by Facebook in 2019 for the Diem blockchain project, emphasizing security, verifiability, and expressiveness. Its innovation lies in embedding asset representation and operations into the type system, thereby preventing common vulnerabilities at the programming language level(blackshear2019move, ). Unlike Solidity on platforms such as Ethereum, Move introduces fundamental solutions to issues like asset duplication, loss, and unauthorized access. As shown in Fig. 1, the designers of Move proposed several core objectives. As shown in the Fig. 2, the core innovation of Move lies in its resource model. A resource is a special data type designed to represent digital assets, such as tokens and NFTs(wang2024smart, ). Inspired by linear logic and linear type systems, this model ensures secure and precise tracking and management of assets and is characterized by several key properties(feist2019slither, ).
Move manages the behavioral characteristics of types through its abilities system, which defines the operations that a type instance can perform. This mechanism is particularly important for resource-type projects. In Move, there are four core abilities. copy, drop, store, and key(li2024scla, ). Resource types typically possess only key and store and lack copy and drop, thereby ensuring that resources cannot be accidentally duplicated or lost. This forms the core of Move’s secure digital asset management. At the storage level, Move employs an address-based global storage model to persist the resources(zou2025malicious, ). Resources with the key ability can be stored in the global state and are associated with specific locations. Move provides several operators, such as move_to<T>, move_from<T>, borrow_global<T>, borrow_global_mut<T>, and exists<T>, to facilitate the management and access of resources in global storage. In light of this, Move’s resource model has created distinctive security design patterns(zhang2022authros, ). The Capability pattern uses resource types as permission tokens, allowing only the code that holds a specific capability to perform privileged operations. The one-time use pattern ensures that certain critical operations can only be executed once, preventing risks from multiple calls. The resource wrapper pattern encapsulates multiple resources within an external resource, enabling systematic and unified management(atzei2017survey, ).
@@ 2.2. Comparison between Move and Solidity
Move differs significantly from Solidity in asset representation, type system, and security mechanisms. While Solidity represents assets via account balance mappings, making them merely part of the contract state and susceptible to duplication, loss, or reentrancy attacks(li2021clue, ), Move treats assets as first-class citizens using resource types. Leveraging a linear type system and the abilities mechanism, Move prevents unintended copying or loss, and its address-based global storage model with resource association enables fine-grained permission control. Move enforces strict error handling via assertions and error codes(larsen2008language, ), and its strong module encapsulation mitigates cross-module attacks and reentrancy. In contrast, Solidity relies on libraries and modifiers for security, which can leave contracts vulnerable to integer overflows, gas DoS, and inconsistent multi-contract state. Performance-wise, MoveVM achieves higher execution efficiency and concurrency, but the ecosystem and tooling are less mature compared to Solidity’s rich development frameworks, libraries, and security audit tools. Overall, Move excels in security and asset management, whereas Solidity offers greater ecosystem maturity and extensibility(hajdu2019solc, ).
@@ 2.3. Move Smart Contract Execution Environment
MoveVM is the virtual machine responsible for executing Move bytecode, handling the loading, verification, and execution of Move programs with a focus on security and resource management(bu2025smartbugbert, ). Its execution workflow includes loading bytecode and input data, verifying the security of bytecode, and initializing the execution environment, which involves setting gas limits and defining the initial state(li2021hybrid, ), executing instructions in sequence, re-verifying execution results, especially the linear availability of resources, and applying state changes or performing rollbacks based on the outcomes of execution. The core components of MoveVM are illustrated in Fig. 3.
@@
The bytecode verifier performs static analysis before deployment and execution to ensure that programs comply with Move’s security requirements, significantly reducing runtime errors and potential vulnerabilities(lahiri2022almost, ). Different blockchain platforms implement Move with varying execution models. Diem employs an account-based storage model, treating multiple operations within a transaction as atomic units and supporting both modules and scripts(li2017discovering, ). Sui introduces an object-based storage model with shared objects, where each resource is an addressable object, supporting parallel transaction execution based on object ownership and multiple-owner access, while simplifying the gas model to improve performance(dwivedi2021legally, ). Aptos uses an enhanced account model that stores resources under accounts, supports parallel transaction execution and Block-STM concurrency control, and provides efficient data structures such as tables, extending the standard library and adding built-in functionalities(tsankov2018securify, ).
Move offers a rich standard library and native functions, typically implemented in a lower-level language, providing operations such as hashing, signature verification, timestamps, randomness, and high-performance computations. These features allow Move programs to interact with external systems while maintaining language simplicity and security(fynn2020smart, ). The gas mechanism controls computational resources and prevents DoS attacks, with each instruction assigned a specific cost. Storage and complex computational operations incur higher costs. Different platforms optimize the gas model(bu2025enhancing, ). Sui uses a simplified gas model to reduce certain operation costs, while Aptos provides finer-grained gas accounting to more accurately reflect resource usage. Understanding and optimizing gas usage is essential for developing efficient Move contracts, reducing transaction costs, and enhancing user experience(thibault2022blockchain, ).
@@ 2.4. Security Threat Analysis of Move Smart Contracts
Move’s resource model is its core security feature, designed to ensure the safe management of digital assets. MoveVM is the virtual machine executing Move bytecode, responsible for loading, verifying, and executing Move programs, with a focus on security and resource management(li2020characterizing, ). Its execution flow includes loading bytecode and input data, verifying bytecode security, initializing the execution environment, including gas limits and initial state, sequentially executing instructions and re-verifying execution results, especially regarding resource linearity, and applying state changes or rolling back depending on the outcome(niu2025natlm, ). The core components of MoveVM include the bytecode verifier, which performs static analysis before deployment and execution to ensure that programs meet Move’s security requirements, significantly reducing runtime errors and potential vulnerabilities(sompolinsky2021phantom, ).
Execution models vary across different blockchain platforms. Diem uses an account-based storage model, treating multiple operations in a transaction as atomic and supporting both modules and scripts. Sui introduces an object-based storage model with shared objects, where each resource is an addressable object, supporting parallel transaction execution based on object ownership and multiple-owner access, and simplifying the gas model to enhance performance(zhou2025blockchain, ). Aptos employs an enhanced account model, storing resources under accounts, supporting parallel transactions and Block-STM concurrency control, while providing efficient data structures such as tables, extending the standard library, and adding built-in functions. Move offers a rich standard library and native functions, often implemented in a low-level language, providing operations such as hashing, signature verification, timestamps, randomness, and high-performance computation(ivanov2023security, ), allowing Move to interact with external systems while maintaining language simplicity and security. The gas mechanism controls computation resources and prevents DoS attacks, with each instruction associated with specific costs, and storage and complex computations incurring relatively high costs. Platforms optimize the gas model differently(jin2025blockchain, ). Sui employs a simplified gas model to reduce some operation costs, while Aptos provides finer-grained gas metering to more accurately reflect actual resource consumption. Understanding and optimizing gas usage is critical for developing efficient Move contracts, reducing transaction costs, and improving user experience(blackshear2020resources, ).
@@ 3. MoveScanner
MoveScanner adopts a modular architecture consisting of three parts. Data processing, analysis engine, and result output. It emphasizes analysis at the bytecode level and can examine compiled Move modules without the need for source code, as illustrated in Fig. 4. Each detection algorithm is independent but shares the underlying control flow and data flow analysis infrastructure to avoid redundant calculations. The overall workflow follows a three-stage model of parsing, analysis, and reporting to ensure scalability and modularity.
@@ 3.1. Core Analysis Technology
MoveScanner implements control flow analysis through the ControlFlowGraph class, which converts bytecode instruction sequences into a graph structure of basic blocks and control edges. Jump instructions are identified as block boundaries, establishing predecessor-successor relationships. The algorithm handles complex nested conditions and loops, which is crucial for detecting resource leaks within conditional branches. Once the control flow graph is constructed, all possible execution paths, including normal and exceptional paths, are represented, providing a foundation for subsequent data flow analysis and resource tracking. Data flow analysis is performed by the DataFlowAnalysis class, including reaching definitions analysis and live variable analysis. Reaching definitions analysis uses a worklist iterative algorithm to track the propagation of variable definitions throughout the program, recording the set of possible definitions at each program point. Live variable analysis operates in a backward direction to determine the set of variables that may be utilized later at each program point. These analyses complement each other. The definition analysis identifies the use of uninitialized variables, while the live variable analysis detects improperly managed resources(brent2018vandal, ).
Cross-module analysis is implemented via the CrossModuleAnalysis class, which constructs a global call graph to detect potential issues in inter-module interactions, as shown in Fig. 5. The capability matrix records access relationships between modules, while the resource flow graph tracks resource propagation across modules. This global perspective allows the system to detect issues that local analysis cannot, such as permission leaks, state contamination, and unauthorized access. Public interface functions are analyzed for parameters and return values to identify potential security boundary violations. Overall, control flow analysis, data flow analysis, and cross-module analysis jointly provide a comprehensive view of variable usage and resource management. For Move’s linear type resources, the system specifically tracks resource creation, movement, and destruction to ensure correct handling along all execution paths(vivar2020analysis, ).
@@ 3.2. Type Systems and Resource Security
MoveScanner conducts a comprehensive type and resource security analysis by first performing full type resolution of Move bytecode via the bytecode.rs module. This module converts SignatureToken in the bytecode into structured ResolvedType, simulates bytecode execution to track stack and local variable states, and provides complete type information and instruction semantics for subsequent vulnerability detection. After parsing, the system constructs a unified intermediate representation (IR) in ir.rs, serving as the foundation for analysis algorithms. The IR adopts a layered structure, with ModuleIR covering module-level information, FunctionIR and StructIR representing functions and structs, and InstructionIR encapsulating instruction-level details. The IR construction involves three key steps. Type resolution, instruction translation, and dependency analysis. Type resolution focuses on moving resource types, mapping raw bytecode type tokens into analyzable forms. Instruction translation simplifies the bytecode instruction set, emphasizing control flow, data flow, and function calls. Dependency analysis generates a module-level call graph to support cross-module security analysis. This layered IR design abstracts the complexity of low-level bytecode, allowing higher-level analysis algorithms to focus on resource security logic, including creation, movement, destruction, and cross-module propagation, thereby enabling precise detection and analysis of potential vulnerabilities in Move smart contracts.

### v3

@@ 1. Introduction
Blockchain technology has evolved continuously since the birth of Bitcoin, from the introduction of smart contracts by Ethereum, and the emergence of third-generation blockchains such as Solana, Aptos, and Sui in recent years. It has gradually transitioned from a single-purpose value-storage system to a programmable distributed application platform. Smart contracts, as a core component of blockchain, grant the system programmability but also introduce new security risks(Luu et al., 2016). The Move language, proposed by the Meta team, was designed with security as its core objective. Its innovative resource model treats digital assets as first-class citizens, achieving non-replicability and non-discardability through a linear type system and explicit move operations, thereby preventing resource leaks and double spending at the language level. Combined with a strict type system, formal verification(Gong et al., 2025), and fine-grained permission controls, Move offers superior security compared to traditional contract languages, such as Ethereum. However, in actual development, Move smart contracts may still contain vulnerabilities due to programming errors, complex business logic, or cross-module interactions(Tsankov et al., 2018).
@@
According to Certik’s 2023 Blockchain Security Report, losses caused by smart contract vulnerabilities exceeded \$2.83 billion, with re-entrancy attacks, permission vulnerabilities, and logical errors as the primary attack types(Liu et al., 2024). The Ethereum ecosystem has developed a relatively mature security analysis toolchain, including Oyente, Securify, ZEUS, Slither, and Mythril, which utilize symbolic execution, abstract interpretation, or static analysis methods to this study lay the foundation for smart contract security research(Li et al., 2025b). In contrast, security tools for the Move ecosystem remain underdeveloped. Move Prover is primarily focused on the formal verification of specific properties, while MoveCheck and VerMove can detect some errors but struggles with complex cross-module vulnerabilities. Manual audit services provided by security companies also lack automation and systematic approaches(Tolmach et al., 2021). This highlights the lack of comprehensive, automated security analysis tools for smart contracts on the Move platform. To address this gap, this paper proposes and implements a static analysis tool, MoveScanne(Bartoletti et al., 2025). The tool constructs a control flow graph and data flow analysis framework, combined with cross-module call tracing technology, enabling the detection of five critical vulnerability categories, including resource leaks, permission defects, and arithmetic overflows(Liu and Li, 2025). MoveScanner adopts a modular architecture, supporting bytecode-level analysis and multi-chain adaptation(Giatzis et al., 2025). They introduced resource trace tracking algorithms and capability matrix analysis methods to reduce the number of false positives. Experimental results show that MoveScanner achieves a detection accuracy of 88.2% in benchmark tests, outperforming existing tools in terms of performance(Kalra et al., 2018).
@@ 2.1. Move Language
Move is a smart contract language specifically designed for the secure management and transfer of digital assets. It was proposed by Facebook in 2019 for the Diem blockchain project, emphasizing security, verifiability, and expressiveness(Song et al., 2024). Its innovation lies in embedding asset representation and operations into the type system, thereby preventing common vulnerabilities at the programming language level(Blackshear et al., 2019). Unlike Solidity on platforms such as Ethereum, Move introduces fundamental solutions to issues like asset duplication, loss, and unauthorized access. As shown in Fig. 1, the designers of Move proposed several core objectives(Karanjai et al., 2024c). As shown in the Fig. 2, the core innovation of Move lies in its resource model. A resource is a special data type designed to represent digital assets, such as tokens and NFTs(Wang et al., 2024c). Inspired by linear logic and linear type systems, this model ensures secure and precise tracking and management of assets and is characterized by several key properties(Feist et al., 2019).
Move manages the behavioral characteristics of types through its abilities system, which defines the operations that a type instance can perform. This mechanism is particularly important for resource-type projects. In Move, there are four core abilities. copy, drop, store, and key(Li et al., 2024). Resource types typically possess only key and store and lack copy and drop, thereby ensuring that resources cannot be accidentally duplicated or lost(Karanjai et al., 2024a). This forms the core of Move’s secure digital asset management. At the storage level, Move employs an address-based global storage model to persist the resources(Zou et al., 2025). Resources with the key ability can be stored in the global state and are associated with specific locations. Move provides several operators, such as move_to<T>, move_from<T>, borrow_global<T>, borrow_global_mut<T>, and exists<T>, to facilitate the management and access of resources in global storage. In light of this, Move’s resource model has created distinctive security design patterns(Zhang et al., 2022). The Capability pattern uses resource types as permission tokens, allowing only the code that holds a specific capability to perform privileged operations(Karanjai et al., 2024b). The one-time use pattern ensures that certain critical operations can only be executed once, preventing risks from multiple calls(Falazi et al., 2024). The resource wrapper pattern encapsulates multiple resources within an external resource, enabling systematic and unified management(Atzei et al., 2017).
@@ 2.2. Comparison between Move and Solidity
Move differs significantly from Solidity in asset representation, type system, and security mechanisms. While Solidity represents assets via account balance mappings, making them merely part of the contract state and susceptible to duplication, loss, or reentrancy attacks(Li et al., 2021), Move treats assets as first-class citizens using resource types. Leveraging a linear type system and the abilities mechanism, Move prevents unintended copying or loss, and its address-based global storage model with resource association enables fine-grained permission control. Move enforces strict error handling via assertions and error codes(Larsen-Freeman and Freeman, 2008), and its strong module encapsulation mitigates cross-module attacks and reentrancy. In contrast, Solidity relies on libraries and modifiers for security, which can leave contracts vulnerable to integer overflows, gas DoS, and inconsistent multi-contract state. Performance-wise, MoveVM achieves higher execution efficiency and concurrency, but the ecosystem and tooling are less mature compared to Solidity’s rich development frameworks, libraries, and security audit tools(Soud et al., 2024). Overall, Move excels in security and asset management, whereas Solidity offers greater ecosystem maturity and extensibility(Hajdu and Jovanović, 2019).
@@ 2.3. Move Smart Contract Execution Environment
MoveVM is the virtual machine responsible for executing Move bytecode, handling the loading, verification, and execution of Move programs with a focus on security and resource management(Bu et al., 2025b). Its execution workflow includes loading bytecode and input data, verifying the security of bytecode, and initializing the execution environment, which involves setting gas limits and defining the initial state(Li and others, 2021), executing instructions in sequence, re-verifying execution results, especially the linear availability of resources, and applying state changes or performing rollbacks based on the outcomes of execution(Wang et al., 2024b). The core components of MoveVM are illustrated in Fig. 3.
@@
The bytecode verifier performs static analysis before deployment and execution to ensure that programs comply with Move’s security requirements, significantly reducing runtime errors and potential vulnerabilities(Lahiri and Roy, 2022). Different blockchain platforms implement Move with varying execution models(Xia et al., 2024). Diem employs an account-based storage model, treating multiple operations within a transaction as atomic units and supporting both modules and scripts(Li et al., 2017). Sui introduces an object-based storage model with shared objects, where each resource is an addressable object, supporting parallel transaction execution based on object ownership and multiple-owner access, while simplifying the gas model to improve performance(Dwivedi et al., 2021). Aptos uses an enhanced account model that stores resources under accounts, supports parallel transaction execution and Block-STM concurrency control, and provides efficient data structures such as tables, extending the standard library and adding built-in functionalities(Tsankov et al., 2018).
Move offers a rich standard library and native functions, typically implemented in a lower-level language, providing operations such as hashing, signature verification, timestamps, randomness, and high-performance computations(Ebekozien et al., 2024). These features allow Move programs to interact with external systems while maintaining language simplicity and security(Fynn et al., 2020). The gas mechanism controls computational resources and prevents DoS attacks, with each instruction assigned a specific cost(Daraghmi et al., 2024). Storage and complex computational operations incur higher costs. Different platforms optimize the gas model(Bu et al., 2025a). Sui uses a simplified gas model to reduce certain operation costs, while Aptos provides finer-grained gas accounting to more accurately reflect resource usage(Zhao et al., 2024). Understanding and optimizing gas usage is essential for developing efficient Move contracts, reducing transaction costs, and enhancing user experience(Thibault et al., 2022).
@@ 2.4. Security Threat Analysis of Move Smart Contracts
Move’s resource model is its core security feature, designed to ensure the safe management of digital assets. MoveVM is the virtual machine executing Move bytecode, responsible for loading, verifying, and executing Move programs, with a focus on security and resource management(Li et al., 2020). Its execution flow includes loading bytecode and input data, verifying bytecode security, initializing the execution environment, including gas limits and initial state, sequentially executing instructions and re-verifying execution results, especially regarding resource linearity, and applying state changes or rolling back depending on the outcome(Niu et al., 2025). The core components of MoveVM include the bytecode verifier, which performs static analysis before deployment and execution to ensure that programs meet Move’s security requirements, significantly reducing runtime errors and potential vulnerabilities(Sompolinsky et al., 2021).
Execution models vary across different blockchain platforms. Diem uses an account-based storage model, treating multiple operations in a transaction as atomic and supporting both modules and scripts. Sui introduces an object-based storage model with shared objects, where each resource is an addressable object, supporting parallel transaction execution based on object ownership and multiple-owner access, and simplifying the gas model to enhance performance(Zhou et al., 2025). Aptos employs an enhanced account model, storing resources under accounts, supporting parallel transactions and Block-STM concurrency control, while providing efficient data structures such as tables, extending the standard library, and adding built-in functions. Move offers a rich standard library and native functions, often implemented in a low-level language, providing operations such as hashing, signature verification, timestamps, randomness, and high-performance computation(Ivanov et al., 2023), allowing Move to interact with external systems while maintaining language simplicity and security. The gas mechanism controls computation resources and prevents DoS attacks, with each instruction associated with specific costs, and storage and complex computations incurring relatively high costs. Platforms optimize the gas model differently(Jin et al., 2025). Sui employs a simplified gas model to reduce some operation costs, while Aptos provides finer-grained gas metering to more accurately reflect actual resource consumption. Understanding and optimizing gas usage is critical for developing efficient Move contracts, reducing transaction costs, and improving user experience(Blackshear et al., 2020).
@@ 3. MoveScanner
MoveScanner adopts a modular architecture consisting of three parts. Data processing, analysis engine, and result output. It emphasizes analysis at the bytecode level and can examine compiled Move modules without the need for source code, as illustrated in Fig. 4. Each detection algorithm is independent but shares the underlying control flow and data flow analysis infrastructure to avoid redundant calculations. The overall workflow follows a three-stage model of parsing, analysis, and reporting to ensure scalability and modularity(Li et al., 2025a).
@@ 3.1. Core Analysis Technology
MoveScanner implements control flow analysis through the ControlFlowGraph class, which converts bytecode instruction sequences into a graph structure of basic blocks and control edges. Jump instructions are identified as block boundaries, establishing predecessor-successor relationships(Karanjai et al., 2025). The algorithm handles complex nested conditions and loops, which is crucial for detecting resource leaks within conditional branches. Once the control flow graph is constructed, all possible execution paths, including normal and exceptional paths, are represented, providing a foundation for subsequent data flow analysis and resource tracking. Data flow analysis is performed by the DataFlowAnalysis class, including reaching definitions analysis and live variable analysis. Reaching definitions analysis uses a worklist iterative algorithm to track the propagation of variable definitions throughout the program, recording the set of possible definitions at each program point. Live variable analysis operates in a backward direction to determine the set of variables that may be utilized later at each program point. These analyses complement each other. The definition analysis identifies the use of uninitialized variables, while the live variable analysis detects improperly managed resources(Brent et al., 2018).
Cross-module analysis is implemented via the CrossModuleAnalysis class, which constructs a global call graph to detect potential issues in inter-module interactions, as shown in Fig. 5. The capability matrix records access relationships between modules, while the resource flow graph tracks resource propagation across modules. This global perspective allows the system to detect issues that local analysis cannot, such as permission leaks, state contamination, and unauthorized access. Public interface functions are analyzed for parameters and return values to identify potential security boundary violations. Overall, control flow analysis, data flow analysis, and cross-module analysis jointly provide a comprehensive view of variable usage and resource management(Hung Lam et al., 2025). For Move’s linear type resources, the system specifically tracks resource creation, movement, and destruction to ensure correct handling along all execution paths(Vivar et al., 2020).
@@ 3.2. Type Systems and Resource Security
MoveScanner conducts a comprehensive type and resource security analysis by first performing full type resolution of Move bytecode via the bytecode.rs module(Chava, 2024). This module converts SignatureToken in the bytecode into structured ResolvedType, simulates bytecode execution to track stack and local variable states, and provides complete type information and instruction semantics for subsequent vulnerability detection(Adam et al., 2024). After parsing, the system constructs a unified intermediate representation (IR) in ir.rs, serving as the foundation for analysis algorithms. The IR adopts a layered structure, with ModuleIR covering module-level information, FunctionIR and StructIR representing functions and structs, and InstructionIR encapsulating instruction-level details. The IR construction involves three key steps. Type resolution, instruction translation, and dependency analysis. Type resolution focuses on moving resource types, mapping raw bytecode type tokens into analyzable forms. Instruction translation simplifies the bytecode instruction set, emphasizing control flow, data flow, and function calls(Adam et al., 2024). Dependency analysis generates a module-level call graph to support cross-module security analysis. This layered IR design abstracts the complexity of low-level bytecode, allowing higher-level analysis algorithms to focus on resource security logic, including creation, movement, destruction, and cross-module propagation, thereby enabling precise detection and analysis of potential vulnerabilities in Move smart contracts(Rao et al., 2024).